# TS-RAG: Retrieval-Augmented Generation based Time Series Foundation Models are Stronger Zero-Shot Forecaster

### v2

@@ 3.2 Retrieve-Augmented Generation based Time Series Foundation Models
By leveraging relevant historical patterns retrieved from an external knowledge database, TS-RAG can enrich the query time series with additional contextual information, thereby improving both model generalization ability and prediction accuracy.
@@
Within TS-RAG, a TSFM has three key components (Tan et al., 2024): an encoding layer, which may include normalization and embedding layers to preprocess and transform the input data; a backbone, typically implemented as a transformer based model (i.e. GPT (Radford et al., 2018), T5 (Raffel et al., 2020), Llama (Touvron et al., 2023b) etc.) to extract temporal representations; and a projection layer, often realized as a multi-layer perceptron (MLP), which maps the temporal representations from the backbone to the final prediction values.
@@
TS-RAG introduces two additional components: a retriever and an augmentation module. These components work alongside the TSFM backbone, enabling the model to adaptively integrate retrieved information and improve forecasting accuracy. More specifically, the encoder from the pretrained Chronos model is used as the embedding model, which generates embeddings for both the query time series context and the contexts stored in the retrieval knowledge database. The retrieval process calculates the Euclidean distance between the query embedding and each stored context embedding in the knowledge base, and then selects the top-$k$ similar candidates based on the smallest distance.
@@
Formally, given a query context $x_{q}$, we first compute its embedding using the Chronos encoder $f_{\text{encoder}}$:
$e_{q}=f_{\text{encoder}}(x_{q}).$
@@
$d(e_{q},e_{i})=\|e_{q}-e_{i}\|_{2},\quad\forall i\in\{1,2,\ldots,n\}.$
To identify the most relevant historical patterns, the retrieval mechanism selects the top-$k$ candidates with the smallest distances:
$\mathcal{C}=\operatorname{TopK}(\{(x_{i},y_{i},d(e_{q},e_{i}))\mid i=1,2,\ldots,n\},k).$
$\operatorname{TopK}(\cdot)$ returns the top-$k$ entries ranked by the lowest distance values. The retrieved set $\mathcal{C}$ contains the most relevant context-forecast pairs, which are subsequently used to augment the forecasting process.
To perform forecasting, we develop a novel Mixture-of-Experts (MoE) augmentation module to integrate the projections of the top-$k$ retrieved forecasting horizons with the query time series embedding from the TSFM backbone to enhance prediction accuracy. Each embedding is treated as an expert, contributing to the final forecast. Initially, each retrieved forecasting horizon $y_{i}$ is encoded independently using a learnable projector:
$\hat{e}_{i}=f_{\text{MLP}}(y_{i}),\quad i=1,2,\dots,k,$
where $f_{\text{MLP}}$ is a feedforward network that maps each retrieved sequence into a dense representation. The resulting embeddings are stacked along a new dimension, forming:
$E_{\text{enc}}=[\hat{e}_{1},\hat{e}_{2},\dots,\hat{e}_{k}]\in\mathbb{R}^{k\times d}$
where $d$ is the embedding dimension. To fuse the retrieved information with the query time series representation $\hat{e}_{q}\in\mathbb{R}^{1\times d}$ generated by the TSFM, the two are concatenated into a single representation:
$E_{\text{concat}}=[\hat{e}_{q};E_{\text{enc}}]\in\mathbb{R}^{(k+1)\times d}.$
This combined representation is passed through a Multi-Head Attention (MHA) layer with residual connection to learn interactions between all the embeddings:
$E_{\text{att}}=\operatorname{MHA}(E_{\text{concat}})+E_{\text{concat}}$
@@
$E_{\text{ffn}}=\operatorname{Dropout}(\operatorname{FFN}(E_{\text{att}}))+E_{\text{att}}.$
A gating mechanism is then applied to adaptively weight the contributions of the retrieved sequences and the model’s original output. Specifically, a gating network computes scores for each expert:
$\alpha=\operatorname{Softmax}(W_{g}E_{\text{ffn}}+b_{g}),$
where $W_{g}$ and $b_{g}$ are learnable parameters, and $\alpha\in\mathbb{R}^{(k+1)\times 1}$ denotes the normalized attention weights. The fused representation is computed as a weighted sum while a skip connection is applied to preserve the TSFM’s original ability:
$e_{\text{final}}=\hat{e}_{q}+\sum_{i=1}^{k+1}\alpha_{i}E_{\text{ffn},i}.$
Finally, the enriched sequence output $e_{\text{final}}$ is passed through the output projection layer of TSFM to generate the final forecast:
$\hat{y}_{q}=f_{\text{proj}}(e_{\text{final}}),$
This mechanism enhances the forecasting in several key aspects. By leveraging retrieved sequences, the model gains access to additional historical information, particularly valuable when the query context alone is insufficient for accurate predictions. The Multi-Head Attention mechanism enables the model to learn context-aware interactions between the retrieved data and its predictions. Next, the gating mechanism adaptively determines the importance of each expert, allowing the model to focus on the most relevant information. Finally, the skip connection ensures that the model’s initial predictions are preserved and enriched, maintaining a balance between internal knowledge and external augmentation. These properties collectively improve the prediction accuracy and enhance the interpretability of the model, particularly in zero-shot forecasting scenarios.
3.3 Adaptive Pretraining and Zero-shot Inference
During the adaptive pretraining phase, we selectively train only the external parameters of the MoE augmentation module in TS-RAG based on pre-constructed multi-domain datasets, while keeping all TSFM parameters frozen.
During zero-shot inference, TS-RAG utilizes its pretrained components to generate forecasts without any task-specific fine-tuning. The retrieval-augmented generation (RAG) approach enables TS-RAG to generalize across diverse forecasting tasks by leveraging external knowledge from a broad set of time series domains. By integrating relevant past trends, the model refines its predictions and mitigates uncertainties, leading to improved forecasting accuracy.
@@ 4.1 Experimental Setup
Datasets and Retrieval Knowledge Base. For the pretraining dataset, we first sample 50 million data points from the Chronos pretraining dataset (Ansari et al., 2024) and further extract a subset of 5 million data points to construct the retrieval knowledge base. To facilitate efficient indexing and retrieval, both the pretraining dataset and the retrieval knowledge base are segmented using a predefined context window. This process results in a total of 26 million pretraining data pairs and 2.8 million retrieval knowledge base pairs.
@@
During the zero-shot inference stage, the historical data of these datasets is expected to contain valuable patterns for augmentation. To utilize this, we build a retrieval knowledge base for each dataset using its own training set. This approach ensures that the retrieval mechanism captures relevant in-domain patterns, enhancing forecasting performance without fine-tuning any parameter.
Baselines. In practice, we use Chronos-Bolt, one of state-of-the-art TSFMs, as the backbone of TS-RAG, as it achieves competitive performance in our evaluations. While TS-RAG is designed to be compatible with any general TSFM, and we have verified its effectiveness beyond Chronos-Bolt, we primarily focus on this backbone in our experiments due to its strong empirical performance. For comparison, we also report the zero-shot performance of other TSFMs, including TTM (Ekambaram et al., 2024), TimesFM (Das et al., 2023), Moirai (Woo et al., 2024), Chronos (Ansari et al., 2024), Chronos-Bolt (Ansari et al., 2024), and MOMENT (Goswami et al., 2024).
Setup. Given that TSFMs are typically trained with a fixed forecasting length (e.g., 64 or 96), we maintain this consistency in both pretraining and zero-shot evaluation. In our setup, the context length is set to 512, and the forecasting length is fixed at 64. We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) as primary evaluation metrics to assess forecasting performance. The detailed definition of the evaluation metrics can be found in Appendix A.4.
Table 1: Long-term zero-shot forecasting results. Best results are highlighted in bold, and second best results are underlined.“—” indicates the datasets were used in pretraining and zero-shot results are not reported.
@@
4.2 Experimental Results for Zero-shot Forecasting
As shown in Table 1, TS-RAGChronos-bolt consistently outperforms other TSFMs, including its backbone model, Chronos-Bolt, across all datasets. This demonstrates the effectiveness of RAG in leveraging relevant time series patterns from the external database to enhance zero-shot forecasting accuracy.
Compared to Chronos-Bolt, TS-RAGChronos-bolt achieves an average reduction of 3.54% in MSE and 1.43% in MAE, confirming that the incorporation of retrieved information improves both precision and robustness. Notably, Chronos-Bolt already demonstrates strong performance on the ETTm1 dataset, achieving an MSE of 0.3109. However, TS-RAGChronos-bolt further reduces MSE by 6.51%, highlighting its ability to refine forecasts even in scenarios where the backbone TSFM is already highly optimized.
Across individual datasets, TS-RAGChronos-bolt consistently achieves the lowest MSE and MAE, demonstrating its robustness across diverse time series patterns. Significant performance gains are observed on datasets such as ETTm1 and Weather, where TS-RAG not only outperforms Chronos-Bolt but also surpasses all other TSFMs by a notable margin. This improvement suggests that RAG is particularly effective in datasets with complex temporal dependencies, where incorporating relevant time series patterns from existing database significantly enhances forecasting accuracy.
@@ 4.3 Ablation Studies
4.3.1 Sensitivity to the Number of Retrieved Sequences
The impact of varying the number of retrieved sequences ($k$) on forecasting performance is illustrated in Figure 2. The x-axis represents the number of retrieved sequences, while the y-axis shows the corresponding Mean Squared Error (MSE). Across all datasets, increasing $k$ initially leads to a significant decrease in MSE, demonstrating that incorporating additional retrieved sequences helps refine predictions by leveraging historical patterns. However, beyond a certain threshold, the improvement plateaus and even decreases slightly in some datasets, indicating diminishing returns as $k$ increases.
Dataset-specific trends further reveal differences in sensitivity to $k$. For instance, ETTm1 and ETTm2 exhibit the most pronounced improvement as $k$ increases, with MSE rapidly declining before stabilizing. This suggests that these datasets benefit significantly from retrieval-augmented learning, likely due to strong temporal dependencies in their historical patterns. ETTh1 and ETTh2 show a similar trend but with a smaller overall reduction in MSE, indicating that while retrieval is beneficial, these datasets may already contain strong intrinsic signals, making additional augmentation less impactful. The Weather, Electricity, and Exchange Rate datasets display a steady decline in MSE with $k$ increasing, but beyond $k=10$, the improvement becomes marginal, suggesting that a moderate number of retrieved sequences is sufficient.
While larger $k$ values generally improve performance, computational costs also increase. Based on the observed results, an optimal range of $k$ between 6 and 10 appears to provide a good trade-off between accuracy and efficiency. These findings suggest that TS-RAG benefits from a carefully chosen number of retrieved sequences rather than an indiscriminate increase.
Figure 2: Parameter sensitivity to the number of retrieved sequences (k) on seven zero-shot evaluation datasets
4.3.2 Effectiveness vs. Different Versions of retrieval knowledge base
The impact of different versions of retrieval knowledge base choices on zero-shot forecasting performance is presented in Table 2. The results compare three settings: (1) w/o RAG, where no retrieval augmentation is applied, (2) pretrained database, where retrieval is performed using a database constructed from the Chronos pretraining dataset, and (3) historical database, where retrieval is performed using a dataset-specific retrieval knowledge base constructed from the training set of each dataset (only for inference, without fine-tuning).
Across two versions of retrieval knowledge bases, TS-RAG consistently outperforms the baseline w/o RAG setting, highlighting the effectiveness of retrieval-augmented generation. Between the two versions of retrieval knowledge bases, the one with the historical database generally yields superior performance, achieving the lowest MSE in six out of seven datasets. This suggests that retrieving in-domain sequences from the same dataset provides more relevant contextual information than retrieving from a broader, pretraining-based retrieval set. However, on the ETTh2 dataset, TS-RAG achieves the best performance when using the pretrained database, implying that retrieval from a more diverse database may sometimes provide a richer set of patterns.
The choice of retrieval knowledge base plays a crucial role in forecasting accuracy. While dataset-specific retrieval tends to be more effective, a multi-domain retrieval knowledge base can sometimes offer advantages, particularly in datasets with high variability. Future work could explore hybrid retrieval mechanisms to further enhance forecasting performance.
Table 2: Long-term zero-shot forecasting results with different retrieval database. The best results are highlighted in bold, and the second-best results are underlined. MSE is reported here.
w/o RAG pretrain database historical database
ETTh1 0.3616 0.3564 0.3557
ETTh2 0.2517 0.2432 0.2451
ETTm1 0.3109 0.2971 0.2906
ETTm2 0.1487 0.1513 0.1466
Weather 0.1525 0.1502 0.1454
Electricity 0.1132 0.1125 0.1120
Exchange rate 0.0673 0.0639 0.0627
4.3.3 Effectiveness of Retrieval Lookback Lengths
Table 3 presents the effect of different retrieval lookback lengths on zero-shot forecasting performance. Given an input sequence of length 512, we explore different retrieval configurations by using only the last 64, 128, or 256 time steps, or the full 512 time steps for retrieval.
Across all datasets, longer retrieval lookback windows (256 or 512) yield relatively better performance, suggesting that incorporating a more extended historical context helps retrieve more relevant sequences. This finding demonstrates that retrieving from longer historical sequences generally improves the quality of retrieved sequences, leading to greater forecasting accuracy. However, in some cases, longer is not always better, indicating that excessive retrieval windows may introduce noise or irrelevant information. This suggests the potential for adaptive retrieval mechanisms that allow the retriever to dynamically determine the most suitable retrieval lookback length for each instance.
w/o RAG 64 128 256 512
ETTh1 0.3616 0.3540 0.3572 0.3539 0.3557
ETTh2 0.2517 0.2432 0.2415 0.2409 0.2451
ETTm1 0.3109 0.3114 0.2935 0.3195 0.2906
ETTm2 0.1487 0.1502 0.1494 0.1518 0.1466
Weather 0.1525 0.1491 0.1526 0.1518 0.1454
Electricity 0.1132 0.1132 0.1125 0.1130 0.1120
Exchange rate 0.0673 0.0678 0.0674 0.0662 0.0627
@@ 4.3.3 Effectiveness of Retrieval Lookback Lengths
Table 4: Zero-shot forecasting results for extended forecasting horizons across multiple datasets. We report MSE.
@@
4.3.4 Effectiveness on Longer Forecasting Horizons
Table 4 presents the zero-shot forecasting results across multiple datasets, including ETTh, ETTm, Weather, Electricity and Exchange Rate. The results show that TS-RAG consistently outperforms its backbone model, demonstrating the effectiveness of retrieval-augmented generation (RAG) in extending prediction horizons while maintaining accuracy. The performance gain suggests that leveraging retrieved sequences mitigates error accumulation, a common issue in rolling-based forecasting.

### v3

@@ 3.2 Architecture of TS-RAG Framework
By leveraging relevant time series contexts retrieved from an external knowledge database, TS-RAG can enrich the input query time series with additional contextual information, thereby improving both model generalization ability and forecasting accuracy.
@@
Within TS-RAG, a TSFM has three key components [47]: an encoding layer, which may include normalization and embedding layers to preprocess and transform the input time series; a backbone, typically implemented as a transformer based model (i.e. GPT [48], T5 [49], Llama [12] etc.) to extract temporal representations; and a projection layer, often implemented as a multi-layer perceptron (MLP), which maps the temporal representations from the backbone to the final prediction values.
@@
TS-RAG introduces two additional components: a Retriever and an Adaptive Retrieval Mixer (ARM) augmentation module. These components work alongside the TSFM backbone, enabling the model to adaptively integrate retrieved information and improve forecasting accuracy. More specifically, the encoder from the pretrained retrieval encoder (e.g., Chronos encoder) is used as the Retriever Encoder, which generates embeddings for both the query time series and the time series contexts stored in the retrieval knowledge database. The Retriever calculates the Euclidean distance between the query embedding and each stored context embedding in the knowledge base, and then selects the top-$k$ similar candidates based on the smallest distance.
@@
Retriever. Formally, given a query context $\mathbf{x}_{q}$, we first obtain its embedding using the retrieval encoder $f_{\text{enc}}$:
$\mathbf{e}_{q}=f_{\text{enc}}(\mathbf{x}_{q}).$ (2)
@@
$d(\mathbf{e}_{q},\mathbf{e}_{i})=\|\mathbf{e}_{q}-\mathbf{e}_{i}\|_{2},\quad\forall i\in\{1,2,\ldots,n\}.$ (3)
To identify the most relevant time series patterns, the retrieval mechanism selects the top-$k$ candidates with the smallest distance:
$\mathcal{C}=\operatorname{TopK}_{\text{min}}\left(\left\{(\mathbf{x}_{i},\mathbf{y}_{i},d(\mathbf{e}_{q},\mathbf{e}_{i}))\mid i=1,2,\ldots,n\right\},k\right),$ (4)
$\operatorname{TopK}_{\text{min}}(\cdot)$ returns the top-$k$ entries ranked by the smallest distance values $d(\mathbf{e}_{q},\mathbf{e}_{i})$. The retrieved set $\mathcal{C}$ contains the most relevant context-forecast pairs, which are subsequently used to augment the forecasting process.
Adaptive Retrieval Mixer (ARM). To perform forecasting, we develop a novel ARM augmentation module to integrate the projections of the top-$k$ retrieved forecasting horizons with the query time series embedding from the TSFM backbone to enhance prediction accuracy. Each embedding is dynamically weighted by the ARM module and contributes accordingly to the final forecast. Initially, each retrieved forecasting horizon $\mathbf{y}_{i}$ is encoded independently using a learnable projector:
$\hat{\mathbf{e}}_{i}=f_{\text{MLP}}(\mathbf{y}_{i}),\quad i=1,2,\dots,k$ (5)
where $f_{\text{MLP}}$ is a feedforward network that maps each retrieved sequence into a dense representation of a $d$-dimensional vector. The resulting embeddings are stacked along a new dimension, forming:
$E_{\text{ret}}=[\hat{\mathbf{e}}_{1},\hat{\mathbf{e}}_{2},\dots,\hat{\mathbf{e}}_{k}]\in\mathbb{R}^{k\times d},$ (6)
where $d$ is the embedding dimension. To fuse the retrieved information with the query time series representation $\hat{\mathbf{e}}_{q}\in\mathbb{R}^{1\times d}$ generated by the TSFM backbone, the two are concatenated into a single representation:
$E_{\text{concat}}=[\hat{\mathbf{e}}_{q};E_{\text{ret}}]\in\mathbb{R}^{(k+1)\times d}.$ (7)
This combined representation is then passed through a Multi-Head Attention (MHA) layer with a residual connection to learn interactions between all the embeddings:
$E_{\text{att}}=\operatorname{MHA}(E_{\text{concat}})+E_{\text{concat}},$ (8)
@@
$E_{\text{ffn}}=\operatorname{Dropout}(\operatorname{FFN}(E_{\text{att}}))+E_{\text{att}}.$ (9)
A mixing mechanism is then applied to adaptively balance the contributions of the retrieved sequence representations and the model’s original representation. Specifically, a scoring network computes an importance weight for each representation:
$\alpha=\operatorname{Softmax}(W_{g}E_{\text{ffn}}+b_{g}).$ (10)
where $W_{g}$ and $b_{g}$ are learnable parameters, and $\alpha\in\mathbb{R}^{(k+1)\times 1}$ denotes the normalized attention weights. The mixed representation is computed as a weighted sum, while a skip connection is applied to preserve the information from the TSFM’s pretrained modeling capability:
$\mathbf{e}_{\text{final}}=\hat{\mathbf{e}}_{q}+\sum_{i=1}^{k+1}\alpha_{i}E_{\text{ffn},i}.$ (11)
Finally, the enriched sequence output $\mathbf{e}_{\text{final}}$ is passed through the output projection layer of TSFM to generate the final forecast:
$\hat{\mathbf{y}}_{q}=f_{\text{proj}}(\mathbf{e}_{\text{final}}),$ (12)
and we follow the same training objective as the TSFM backbone (see Setup in Section 4.1).
The ARM mechanism enhances forecasting in several key aspects. By leveraging retrieved sequences, the model gains access to additional information, which is particularly valuable when the query context alone is insufficient for accurate predictions. The Multi-Head Attention mechanism enables the model to learn context-aware interactions between the retrieved data and its predictions. Next, the mixing mechanism adaptively determines the importance of each candidate, allowing the model to focus on the most relevant information. Finally, the skip connection ensures that the model’s initial predictions are preserved and enriched, maintaining a balance between query knowledge and external augmentation. These designs collectively improve the prediction accuracy and enhance the interpretability of the model, particularly in zero-shot forecasting scenarios.
3.3 Pretraining Strategy and Zero-shot Inference
Pretraining Strategy. For pretraining, we selectively only train the external parameters of the projector and the ARM augmentation module in TS-RAG based on pre-constructed multi-domain datasets, while keeping all other parameters (the TSFM backbone and the Retrieval Encoder) frozen.
Zero-shot Inference. During the zero-shot inference stage, TS-RAG utilizes its pretrained components to generate forecasts without any task-specific fine-tuning. The RAG approach enables TS-RAG to generalize across diverse forecasting tasks by leveraging external knowledge from a broad set of time series domains. Our experiments in Section 4.3 demonstrate the effectiveness of TS-RAG in in-domain, distribution shift, cross-domain, and multi-domain settings.
@@ 4.1 Experimental Setup
Datasets and Retrieval Knowledge Base. For the pretraining dataset, we first uniformly sample 50 million data points from the Chronos pretraining dataset [25] and further uniformly sample a subset of 5 million data points to construct the multi-domain retrieval knowledge base. To facilitate efficient indexing and retrieval, both the pretraining dataset and the retrieval knowledge base are segmented using a predefined context window. This process results in a total of 26 million pretraining data pairs and 2.8 million retrieval knowledge base pairs.
@@
During zero-shot inference, the retrieval knowledge base can be constructed in various ways, including in-domain, distribution shift, cross-domain, and multi-domain settings. We further discuss the setup and impact of different knowledge base choices in Section 4.3.
Baselines. In practice, we use Chronos-Bolt, one of the state-of-the-art TSFMs, as the backbone of TS-RAG, as it achieves competitive performance in our evaluations. TS-RAG is designed to be compatible with any general TSFM. While our main experiments primarily use Chronos-Bolt as the backbone due to its strong empirical performance, we also verify TS-RAG’s effectiveness with MOMENT [37] in Appendix B.1. For comparison, we also report the zero-shot performance of other TSFMs, including TTM [24], TimesFM [23], Moirai [26], Chronos [25], Chronos-Bolt [25], MOMENT [37], and Time-MoE [41].
Setup. Given that TSFMs are typically trained with a fixed forecasting length (e.g., 64 or 96), we maintain this consistency in both pretraining and zero-shot evaluation. The context length is set to 512, and the forecasting length is fixed at 64. We adopt the same forecasting loss as the backbone TSFM; when using Chronos-bolt as the backbone, we apply the quantile regression loss following its original implementation. Mean Squared Error (MSE) and Mean Absolute Error (MAE) are used as primary evaluation metrics, with detailed definitions provided in Appendix A.5.
4.2 Experimental Results for Zero-shot Forecasting
As shown in Table 1, TS-RAGChronos-bolt consistently outperforms other TSFMs, including its backbone Chronos-Bolt, across all datasets, demonstrating its effectiveness in leveraging external patterns to enhance zero-shot forecasting.
Compared to Chronos-Bolt, TS-RAGChronos-bolt achieves an average reduction of 3.54% in MSE and 1.43% in MAE, confirming that the incorporation of retrieved information improves both precision and robustness. Notably, Chronos-Bolt already performs well on the Exchange Rate dataset, achieving an MSE of 0.0673, yet TS-RAGChronos-bolt further reduces the MSE by 6.84%, demonstrating its ability to refine forecasts even when the backbone is highly optimized.
Across each individual dataset, TS-RAGChronos-bolt consistently achieves the lowest MSE and MAE, demonstrating its robustness across diverse time series patterns. Significant performance gains are observed on datasets such as ETTm1 and Weather, where TS-RAG not only outperforms Chronos-Bolt but also surpasses all other TSFMs by a notable margin. This improvement suggests that RAG is particularly effective in datasets with complex temporal dependencies, where incorporating relevant time series patterns from an existing database significantly enhances forecasting accuracy.
Table 1: Long-term zero-shot forecasting results. Best results are highlighted in bold, and second best results are underlined.“—” indicates the datasets were used in pretraining and zero-shot results are not reported. More results are in Appendix B.2 Table 5.
@@ 4.3 Ablation Studies
Effectiveness of ARM Module
The augmentation module plays a critical role in TS-RAG, as it determines how the retrieved forecasting horizons are integrated with the query representation. An ineffective fusion design could limit the benefits of retrieval, leading to sub-optimal forecasting performance. To assess the contribution of the ARM module, we conduct ablation experiments where we replace the ARM with a simpler alternative: a gated fusion module that directly linearly combines the TSFM’s forecast with the retrieved forecasting horizon.
We compare the two augmentation modules under the same pretrain-zeroshot setting, keeping all other components fixed. Full results are in Appendix B.2 Table 7. As shown in Figure 2 (left), although the gated fusion variant ("W/Gate") also brings performance improvements over the TSFM baseline, the gains are consistently lower than those achieved by the ARM-based TS-RAG ("W/ARM"). This result highlights the effectiveness of the ARM module in dynamically mixing the contributions of different retrieved patterns as well as the original output representation. The attention-based fusion with residual connection in ARM enables richer interactions and adaptive integration, which proves to be crucial for zero-shot forecasting.
Impact of Retrieval Knowledge Base
We evaluate the impact of four retrieval knowledge base configurations on zero-shot forecasting performance, as shown in Figure 2 (right). The configurations include: (1) in-domain, the retrieval knowledge base is built using the training set of the same dataset; (2) distribution shift, the retrieval knowledge base is constructed using the training set from a closely related dataset with a different distribution (e.g., using the ETTh2 training set when evaluating on ETTh1); (3) cross-domain, using different domain data; and (4) multi-domain, using data from multiple domains (as mentioned in the experimental setup, we construct the multi-domain knowledge base using subsets of the pretraining dataset).
The detailed results are in Appendix B.2 Table 6. Across all configurations, TS-RAG improves forecasting performance over the baseline, particularly in MSE, where in-domain retrieval consistently achieves the lowest error. These results suggest that while retrieval generally enhances forecasting, the composition of the retrieval knowledge base plays a crucial role.
Figure 2: Comparison of average MSE and MAE across augmentation methods (left) and knowledge bases (right). w/o RAG indicates Chronos-bolt baseline; w/ ARM and w/ Gate refer to TS-RAG with ARM and Gate augmentation modules, respectively. More detailed results are in Appendix B.2
Effectiveness of Longer Forecasting Horizons
Table 2 presents the zero-shot forecasting results across multiple datasets, including ETTh, ETTm, Weather, Electricity, and Exchange Rate. The results show that TS-RAG consistently outperforms its backbone model, demonstrating the effectiveness of RAG in extending prediction horizons while maintaining accuracy. The performance gain suggests that leveraging retrieved sequences mitigates error accumulation, a common issue in rolling-based forecasting.
@@
Table 2: Zero-shot forecasting results for extended forecasting horizons across multiple datasets. We report MSE. More results are in Appendix B.2 Table 8.
@@
Additional Ablation Study
We further conduct additional ablation studies to investigate the impact of retriever configurations, including the number of retrieved sequences, retrieval lookback length, and retriever encoder choice. The results confirm that while TS-RAG benefits from longer lookback windows and a moderate number of retrieved sequences, it is robust to the choice of retriever encoder. Detailed results and analysis are provided in Appendices B.3, B.4, and B.5.
4.4 Comparison with RAF
Table 3: Zero-shot forecasting comparison between TS-RAG and RAF. (Left) Average MSE and MAE for 512-64 forecasting across seven benchmarks. (Right) Average Inference Speed on ETTh datasets. More results are in Appendix B.6.
Method Average MSE Average MAE
RAF 0.2318 0.2738
TS-RAG (Ours) 0.1940 0.2492
Method Retrieval Forward Pass Total
RAF 3290 ms/iter 184 ms/iter 3474 ms/iter
TS-RAG (Ours) 9.2 ms/iter 0.44 ms/iter 9.62 ms/iter
To further evaluate the performance of TS-RAG, we compare it with RAF (Retrieval Augmented Forecasting) [46], a recent retrieval-augmented method for zero-shot time series forecasting. We conduct the comparison from two aspects: effectiveness and efficiency. More implementation details can be found in Appendix A.3.
Effectiveness
We evaluate the zero-shot forecasting performance of TS-RAG and RAF on seven benchmark datasets. TS-RAG achieves an average MSE of 0.1940, outperforming RAF, which records an average MSE of 0.2320. Moreover, TS-RAG consistently delivers lower errors across all datasets, highlighting its superior generalization capability and robustness.
Efficiency
We also compare the inference speed of TS-RAG and RAF on the ETTh dataset. TS-RAG significantly outperforms RAF in terms of inference efficiency. Specifically, for the retrieval stage, TS-RAG completes each iteration in just 9.2 ms, whereas RAF requires 3290 ms per iteration. This substantial speed improvement mainly comes from two factors: (1) the use of FAISS for fast nearest-neighbor search, and (2) TS-RAG conducts retrieval over compact embeddings.
4.5 Interpretable Forecasting with Case Studies
TSFMs are often used as black boxes, making it difficult to understand how predictions are made. TS-RAG addresses this by providing two key interpretability features: (1) retrieval-as-evidence, which surfaces top-$k$ analogue sequences for each query window, and (2) transparent weighting, which highlights the most relevant subset of retrieved sequences that have the greatest influence through similarity scores.
We illustrate these capabilities with case studies on retrieval quality and its impact on forecasting performance. Figure 4 shows the retrieval process on the Weather dataset, where the retriever selects highly relevant sequences aligned with both trend and periodicity of the query, serving as explicit evidence for the forecast. Figure 4 presents a case from the ETTm1 dataset, where TS-RAG improves forecast accuracy by leveraging retrieved sequences that exhibit similar sharp downward trends, which the backbone TSFM fails to capture. The retrieved forecasting horizon shown corresponds to the retrieved sequence with the highest weighting, highlighting the model’s ability to transparently surface and utilize the most influential retrieved patterns.
These examples demonstrate that TS-RAG not only improves accuracy but also makes its forecasts more interpretable by exposing the key retrieved patterns and their contributions. More case studies are provided in Appendix C.